# SGLC: Semantic Graph-Guided Coarse-Fine-Refine Full Loop Closing for LiDAR SLAM

### v1

@@ IV-A Experimental Setup
Dataset. We evaluate our method on two publicly available datasets: KITTI [15] and KITTI-360 [16]. KITTI includes 11 sequences with ground truth poses, among which six sequences (00, 02, 05, 06, 07, 08) contain loop closures and sequence 08 contains reverse loop closures. KITTI-360 is composed of 9 sequences, with 6 of them containing loops. Compared to KITTI, KITTI-360 has a greater number of loop closures and reverse loop closures.
Implement Details. In our experiments, we set $d_{\text{max}}=60$, $k=30$, and the graph and background similarity thresholds are $0.58$ and $0.7$, respectively. We use the semantic information from a semantic segmentation network [25] on both datasets and mainly build semantic graphs from $static\;vehicle$, $pole$ and $trunk$ instances, and utilize background points from $building$, $fence$, $road$ and $vegetation$. This is convenient for replacing or adding other semantic information. All the experiments are conducted on a machine with an AMD 3960X @3.8GHz CPU and a NVIDIA RTX 3090 GPU. More implementation details can be found in our open-source repository.
TABLE II: The performance evaluation of loop closure detection on the KITTI dataset using overlap-based loop pairs.
@@
Handcrafted Scan Context 0.836 0.835 0.820 0.869
LiDAR-Iris 0.843 0.848 0.835 0.877
GOSMatch 0.926 0.844 0.945 1.0
SSC 0.937 0.913 0.906 0.959
@@
Ours 0.953 0.934 0.957 0.989
  • The best results are highlighted in bold and the second results are underlined.
TABLE III: The comparison of loop pose estimation errors on the KITTI dataset
@@
RR(%) RTE(m) RYE() RR(%) RTE(m) RYE() RR(%) RTE(m) RYE() RR(%) RTE(m) RYE() Time(ms)
BoW3D 95.50 0.06 0.84 77.36 0.09 2.01 57.02 0.07 0.96 47.21 0.10 1.93 118.1
LCDNet(fast) 93.03 0.65 0.86 60.71 1.02 1.65 66.82 0.56 0.68 47.29 0.88 1.06 229.3
LCDNet 100 0.11 0.12 100 0.15 0.34 94.86 0.23 0.24 96.87 0.31 0.51 576.7
Ours 100 0.04 0.18 100 0.07 0.39 99.94 0.04 0.19 99.60 0.07 0.40 97.3
  • denotes the entire runtime including the semantic segmentation. The best results are highlighted in bold and the second results are underlined.
TABLE IV: The comparison of loop pose estimation errors on the KITTI-360 dataset.
Seq.0002 Seq.0009 Time
RR(%) RTE(m) RYE() RR(%) RTE(m) RYE() (ms)
BoW3D 67.62 0.19 2.03 87.40 0.13 1.27 120.8
LCDNet(fast) 82.92 0.84 1.28 89.49 0.76 0.99 254.4
LCDNet 98.62 0.28 0.32 100 0.18 0.20 627.9
Ours 95.14 0.20 0.85 99.87 0.11 0.48 98.9
  • denotes the entire runtime including the semantic segmentation. The best results are highlighted in bold and the second results are underlined.
IV-B Loop Closure Detection
We follow the experimental setups of Li et al. [10] to evaluate loop closure detection performance. For the KITTI dataset, we perform performance comparisons on all sequences with loop closure. For the KITTI-360 dataset, in line with [11], we focus on evaluation sequences 0002 and 0009 with the highest number of loop closures. In the experiments, we regard LiDAR scan pairs as positive samples of loop closure when their Euclidean distance is less than 3 m and as negative samples if the distance exceeds 20 m.
Metrics. Following [10, 12], we use the maximum $F_{1}$ score and Extended Precision ($EP$) as evaluation metric:
$F_{1}=2\times\frac{P\times R}{P+R}$ (9)
$EP=\frac{1}{2}(P_{R0}+R_{P100})$ (10)
where $P,R$ denote the precision and recall, respectively. $P_{R0}$ is the precision at minimum recall and $R_{P100}$ the max recall at 100% precision.
Baselines. We compare the results with SOTA baseline methods, including DNN-based methods: PointNetVLAD [2], OverlapNet [9], LCDNet [11], OverlapTransformer (OT) [5], BEVPlace [6], as well as handcrafted-based methods: Scan Context [7], LiDAR-Iris [8], GOSMatch [14], SSC [10], BoW3D [12].
Results. The results are shown in Tab. I. Our method outperforms SOTA on multiple sequences and achieves the best average F1max score and $EP$ for the KITTI dataset. Specifically, for sequence 08 which only contains reverse loop closure, our approach still exhibits superior performance while most handcrafted-based methods suffer from significant degradation. This proves that our method possesses good rotational invariance. For the KITTI-360 dataset, our method can still get competitive results, demonstrating its generalization capability.
Furthermore, to investigate the loop closure detection performance at further distances, we followed [9, 5] to regard two LiDAR scans as a loop closure when their overlap ratio is beyond 0.3, which indicates that the maximum possible distance between them is around 15 m. We adopt the same experimental setup as theirs to evaluate our method on the KITTI 00 sequence using AUC, F1max, Recall@1, and Recall@1% as metrics. The results are shown in Tab. II. Due to BoW3D leaning more towards geometric verification, we are unable to generate its AUC and recall@1% results from its open-source implementation. From the results, our method significantly outperforms SOTA baselines in terms of overall metric, indicating its robustness for detecting long-distance loop closures. Although GOSMatch also has a commendable recall capability, its F1max score is slightly inferior to other methods, indicating an insufficient capability to accurately determine true loop closure.
Figure 3: The qualitative comparison of loop pose estimation on the KITTI dataset using overlap-based loop pairs. Dashed ellipses are directly annotated on the registration results, while solid boxes indicate local magnification.
@@ IV-C Loop Pose Estimation
To validate the loop pose estimation performance, we follow Cattaneo et al. [11] and evaluate our approach on the sequences 00 and 08 of the KITTI dataset, sequences 0002 and 0009 of the KITTI-360 dataset. To keep consistent with them, we choose the loop closure samples when the distance is within 4 m. Additionally, we also select the more challenging closed-loop pairs with low overlap for testing.
Metric. We apply three evaluation metrics for comparisons: (i) Registration Recall (RR): represents the percentage of successful alignment. (ii) Relative Translation Error (RTE), the Euclidean distance between the ground truth and estimated poses. (iii) Relative Yaw Error (RYE), the yaw angle difference between the ground truth and estimated poses. Two scans are regarded as a successful registration when the RTE and RYE are below 2 m and 5, respectively.
@@
Results. As shown on the left of Tab. III, our method achieves competitive results compared to LCDNet, and exceeds LCDNet (fast) and BoW3D by a large margin on the KITTI dataset of the closed loop within 4 meters. Both in the same direction loop correction in sequence 00 and reverse loop correction in sequence 08, our method and LCDNet get a high registration recall (RR) while BoW3D and LCDNet (fast) degrade significantly in reverse loop correction. In terms of RTE, our method excels over LCDNet, demonstrating its superior performance in translation estimation. To further highlight the gap between these methods, we test them on loop pairs with an overlap ratio exceeding 0.3. The comparison results are presented on the right of Tab. III. It can be seen that our method still maintains strong registration performance and surpasses all others across all metrics, emphasizing its robustness. In comparison, LCDNet experienced a performance decline when performing registration with low overlap. We also present the qualitative results in Fig. 3. As can be seen, our method achieves better alignment even in low overlap loop closure, attributing to the semantically-guided coarse-fine-refine point cloud alignment process. Tab. IV shows the results on the KITTI-360 dataset with close loop within 4 m. We also report the average running time for all methods on our machine. From the registration results, our method remains competitive compared to other baselines, demonstrating the generalizability of our approach. Additionally, our method achieves the fastest running speed and significantly outperforms other methods even when the time of semantic segmentation is included, approximately 67.1 ms for SegNet4D [25]. By comparison, LCDNet operates with extreme time consumption, rendering it unsuitable for real-time SLAM systems.
@@
IV-D Performance on SLAM System
We integrate our method into the A-LOAM odometry to eliminate cumulative errors for evaluating the accuracy of SLAM trajectories, and compare the results with baseline integrated with BoW3D. We utilize a Incremental Smoothing and Mapping (iSAM2) [32] based pose-graph optimization (PGO) [^1] https://github.com/gisbi-kim/SC-A-LOAM framework to build a factor graph for managing the the keyframe poses. If a loop closure is detected, we will add this loop closure constraint into the factor graph to eliminate errors and ensure global consistency. As shown in Fig. 4, we present the trajectory error analysis between keyframe and ground truth poses. From the results, our method significantly enhances the SLAM system by reducing the cumulative odometry error, and shows superior performance compared to the baseline that incorporates BoW3D. This improvement is attributed to the excellent loop closing capability of our approach.
TABLE V: Ablation studies on KITTI 08 sequence. “Node Glo.” denotes the global properties of instance node, mentioned in Sec. III-A. “Out. Pru.” denotes outlier pruning for node correspondences. “Den. ICP” denotes dense points registration, i.e., $\mathbf{T}_{\text{icp}}$. “P-P Ali.” denotes point-plane alignment, i.e., $\mathbf{T}_{\text{refine}}$.
Node Glo. Out. Pru. Den. ICP P-P Ali. F1max RR(%) RTE(m) RYE() Time (ms)
[[[[A]]]] 0.981 97.20 0.23 1.31 10.2
[[[[B]]]] 0.987 97.80 0.22 1.29 +0.6
[[[[C]]]] 0.988 98.61 0.20 0.98 +1.4
[[[[D]]]] - 99.97 0.08 0.50 +13.2
[[[[E]]]] - 99.53 0.08 0.42 +7.9
[[[[F]]]] - 100 0.07 0.39 +21.8
  • “+” denotes the time increment compared to [[[[A]]]]. The best results are highlighted in bold
@@ IV-E Ablation Study
In this section, we conduct a series of ablation studies to evaluate the effectiveness of designed components and analyze their runtime. The results are presented in Tab. V. The Experiment [[[[A]]]] is our baseline by removing those components listed in Tab. V. Comparing Experiment [[[[A]]]] and [[[[B]]]], it is demonstrated that incorporating the global properties of instance nodes can enhance the performance of loop closing. As anticipated, it increases each node’s distinctiveness, which is beneficial for finding the correct node correspondences. In Experiment [[[[C]]]], we can see the effectiveness of outlier pruning in improving registration accuracy with the same number of RANSAC iterations as [[[[B]]]]. For Experiment [[[[D]]]], [[[[E]]]] and [[[[F]]]], dense points registration and point-plane alignment are applied mainly for loop scan determined after geometric verification, so they do not affect the results of loop closure detection. The results clearly show that both dense points registration and point-plane alignment can improve registration performance, and we can get optimal results only when they operate in conjunction. Besides, each component in our framework exhibits high execution efficiency from the runtime perspective.

### v2

@@ IV-A Experimental Setup
Dataset. We evaluate our method on KITTI [16] and KITTI-360 [17] datasets. Additionally, Ford Campus [18] and Apollo [19] datasets are used to test its generalization ability.
Implementation Details. The parameters used in our experiments are listed in Tab. I. Due to background similarity being easily affected by semantic segmentation quality, we set $t_{b}$ for different datasets ($t_{b}=0.5$ for KITTI, KITTI-360 and Apollo, $t_{b}=0.45$ for Ford Campus). Other parameters remain unchanged. We use the semantic information from a semantic segmentation network [30] on all datasets and we train different models on the KITTI dataset to ensure that each test sequence remains unseen during the training phase. We mainly build semantic graphs from $static\;vehicle$, $pole$ and $trunk$ instances, and utilize background points from $building$, $fence$, $road$ and $vegetation$. This is convenient for replacing or adding other semantic information. To fairly compare all semantic-assisted loop closing methods, we evaluate them using the same labels from SegNet4D [30]. All the experiments are conducted on a machine with an AMD 3960X @3.8GHz CPU and a NVIDIA RTX 3090 GPU. More implementation details can be found in our online supplementary  [^1] https://github.com/nubot-nudt/SGLC/Supplementary.pdf.
TABLE I: The parameters settings of SGLC.
Parameter Value Description
$d_{\text{max}}$ 60m edge threshold for semantic graph
$d_{i}$ 2m interval for edge histogram descriptors
$k$ 30 the dimension for eigenvectors
$d_{b}$ 2m difference threshold of bounding box dimension
$d_{l}$ 20m the range for generating local triangle
$t_{m}$ 2 threshold for determining true node correspondences
$t_{g}$ 0.5 threshold for semantic graph similarity
$t_{b}$ 0.5 threshold for background similarity
$D$ 120 the dimension for node descriptor
$D^{\prime}$ 231 the dimension for LiDAR scan descriptor
TABLE II: The performance comparison of F1 max scores and Extended Precision on the KITTI and KITTI-360 datasets.
KITTI KITTI-360 Time
Methods 00 02 05 06 07 08 Mean 0002 0009 (ms)
DNN-based PV 0.779/0.641 0.727/0.691 0.541/0.536 0.852/0.767 0.631/0.591 0.037/0.500 0.595/0.621 0.349/0.515 0.330/0.510 11.3
SGPR 0.720/0.507 0.823/0.531 0.720/0.552 0.680/0.524 0.700/0.500 0.683/0.506 0.721/0.520 0.862/0.506 0.845/0.503 8.8
LCDNet 0.970/0.847 0.966/0.917 0.969/0.938 0.958/0.920 0.916/0.684 0.989/0.908 0.961/0.869 0.998/0.993 0.988/0.734 147.7
OT 0.873/0.800 0.810/0.725 0.837/0.772 0.876/0.809 0.625/0.505 0.667/0.510 0.781/0.687 0.796/0.507 0.879/0.528 22.8
BEVPlace 0.960/0.849 0.845/0.819 0.885/0.815 0.895/0.815 0.917/0.687 0.967/0.868 0.912/0.809 0.920/0.504 0.938/0.684 33.9
PADLoC 0.983/0.912 0.920/0.880 0.957/0.863 0.981/0.944 0.781/0.704 0.910/0.718 0.922/0.837 0.948/0.596 0.969/0.797 221.5
Handcrafted SC 0.750/0.609 0.782/0.632 0.895/0.797 0.968/0.924 0.662/0.554 0.607/0.569 0.777/0.681 0.771/0.554 0.851/0.619 13.9
GOSMatch 0.916/0.535 0.694/0.575 0.785/0.611 0.491/0.518 0.947/0.913 0.908/0.571 0.790/0.621 0.694/0.500 0.766/0.575 6.4
SSC 0.955/0.865 0.933/0.768 0.966/0.956 0.983/0.978 0.894/0.785 0.950/0.940 0.947/0.882 0.965/0.793 0.966/0.873 4.6
BoW3D 0.977/0.981 0.578/0.704 0.965/0.969 0.985/0.985 0.906/0.929 0.900/0.866 0.885/0.906 0.210/0.560 0.682/0.761 78.0
CC 0.977/0.964 0.903/0.568 0.958/0.901 0.993/0.959 0.906/0.811 0.823/0.575 0.927/0.796 0.717/0.560 0.886/0.620 8.4
Ours 0.998/0.986 0.888/0.899 0.969/0.967 0.995/0.963 0.993/0.991 0.988/0.980 0.972/0.964 0.932/0.934 0.994/0.978 10.2
  • [ $F_{1}$ max / $EP$ ], the best results are highlighted in bold and the second best are underlined. denotes semantic-assisted method.
TABLE III: The performance evaluation of loop closure detection on the KITTI dataset using overlap-based loop pairs.
@@
PADLoC 0.934 0.903 0.930 0.975
Handcrafted SC 0.836 0.835 0.820 0.869
GOSMatch 0.906 0.829 0.941 0.997
SSC 0.924 0.882 0.900 0.951
@@
CC 0.873 0.902 0.865 0.868
Ours 0.949 0.931 0.950 0.986
IV-B Loop Closure Detection
We follow the experimental setups of Li et al. [10] to evaluate LCD performance and regard LiDAR scan pairs as positive samples of loop closure when their Euclidean distance is less than 3 m, as negative samples if the distance exceeds 20 m. For the KITTI dataset, we perform performance comparisons on all sequences with loop closure. For the KITTI-360 dataset, in line with [12], we focus on sequences 0002 and 0009 with the highest number of loop closures. Besides, we also report the results on the Ford Campus dataset (sequence 00) and a subset of the Apollo Columbia Park MapData proposed by [37] for generalization evaluation.
Metrics. Following [10, 13], we use the maximum $F_{1}$ score and Extended Precision ($EP$) as evaluation metric.
Baselines. We compare the results with SOTA baselines, including DNN-based methods: PointNetVLAD (PV) [2], SGPR [3], LCDNet [12], OverlapTransformer (OT) [5], BEVPlace [6], PADLoC [29], as well as handcrafted-based methods: Scan Context (SC) [7], GOSMatch [15], SSC [10], BoW3D [13], Contour Context (CC) [11].
Results. The results are shown in Tab. II. Our method outperforms SOTA on multiple sequences and achieves the best average F1max score and $EP$ for the KITTI dataset. Specifically, for sequence 08 with many reverse loop closures, our approach still exhibits superior performance, proving its good rotational invariance. And on the KITTI360 dataset, our method remains competitive. Additionally, we report the average runtime for all methods, including descriptor generation and loop closure determination. Our method performs as fast as handcrafted methods while achieving superior performance in most cases, demonstrating both efficiency and effectiveness.
Furthermore, to investigate the loop closure detection performance at further distances, we followed [9, 5] regarding two LiDAR scans as a loop closure when their overlap ratio is beyond 0.3, which indicates that the maximum possible distance between them is around 15 m. And we adopt the same experimental setup as theirs to evaluate our method on the KITTI 00 sequence using AUC, F1max, Recall@1, and Recall@1% as metrics. The results are shown in Tab. III. Due to BoW3D leaning more towards geometric verification, we are unable to generate its AUC and recall@1% results from its open-source implementation. From the results, our method significantly outperforms SOTA baselines in terms of overall metric, indicating its robustness.
To evaluate the generalization capability of SGLC, we conducted additional experiments on the Ford Campus and Apollo datasets, focusing on loop pairs with an overlap ratio greater than 0.3. We primarily compared semantic-assisted methods to assess their adaptability to scene semantic changes or label quality declines. Due to the page limitation, we show the semantic segmentation details in our online supplementary. As shown in the Tab. IV, our method demonstrates the best performance on both the Ford Campus and Apollo datasets, highlighting its strong generalization capability. When the quality of semantic segmentation declines, some instances may be partially segmented, but after clustering, they still form a single object, which does not affect the construction of our semantic graph. Once the semantic graph is built, the LiDAR scan descriptors retrieve multiple candidate scans, including the correct loop scan. Through our geometric verification via graph matching, the true loop closure is identified. Besides, for short-term loop closing in SLAM, static vehicles can serve as valuable feature nodes to enhance loop detection.
TABLE IV: The generalization performance evaluation.
Ford Campus Apollo
Methods AUC F1max Recall @1 Recall @1% AUC F1max Recall @1 Recall @1%
SGPR 0.412 0.439 0.467 0.951 0.640 0.451 0.626 0.936
GOSMatch 0.752 0.632 0.820 0.954 0.677 0.568 0.739 0.934
SSC 0.924 0.865 0.915 0.964 0.937 0.916 0.917 0.943
PADLoC 0.938 0.873 0.910 0.963 0.721 0.609 0.735 0.910
Ours 0.955 0.908 0.920 0.970 0.956 0.910 0.933 0.951
Ours 0.959 0.897 0.896 0.960 0.957 0.919 0.956 0.974
  • denotes build semantic graph without vehicle node. The best results are highlighted in bold and the second best are underlined.
TABLE V: The comparison of loop pose estimation errors on the KITTI dataset
@@ IV-B Loop Closure Detection
RR(%) RTE(m) RRE() RR(%) RTE(m) RRE() RR(%) RTE(m) RRE() RR(%) RTE(m) RRE() Time(ms)
BoW3D 95.61 0.06 0.81 77.29 0.09 2.06 57.69 0.07 0.92 46.10 0.10 1.95 40.0
LCDNet(fast) 97.44 0.52 0.60 78.28 0.99 1.29 66.70 0.56 1.02 47.20 0.88 1.29 81.7
LCDNet 100 0.13 0.44 100 0.20 0.57 93.51 0.21 0.81 95.39 0.31 0.94 429.1
Ours 100 0.04 0.21 100 0.08 0.41 99.82 0.04 0.24 99.57 0.08 0.46 20.8
  • The best results are highlighted in bold and the second best are underlined.
TABLE VI: Loop pose estimation results on the KITTI-360 dataset.
Seq.0002 Seq.0009
RR(%) RTE(m) RRE() RR(%) RTE(m) RRE()
BoW3D 67.73 0.20 1.99 87.32 0.13 1.31
LCDNet(fast) 82.92 0.84 1.17 98.81 0.42 1.07
LCDNet 98.56 0.28 1.03 100 0.19 0.74
Ours 95.01 0.21 0.89 99.40 0.11 0.51
@@ IV-C Loop Pose Estimation
To validate the loop pose estimation performance, we follow Cattaneo et al. [12] and evaluate our approach on the KITTI sequences 00 and 08 and KITTI-360 sequences 0002 and 0009. To keep consistent with them, we choose the loop closure samples when the distance is within 4 m. Additionally, we also select the more challenging loop pairs with overlap rate beyond 0.3 for testing.
Metric. (i) Registration Recall (RR), represents the percentage of successful alignment. (ii) Relative Translation Error (RTE), the Euclidean distance between the ground truth and estimated poses. (iii) Relative Rotation Error (RRE), the rotation angle difference between the ground truth and estimated poses. Two scans are regarded as a successful registration when the RTE and RRE are below 2 m and 5, respectively.
@@
Results. As shown on the Tab. V, Our method achieves the best loop pose estimation performance on both distance-based loop pairs and more challenging overlap-based loop pairs. Initially, both LCDNet and our method achieve a 100% RR for distance-based loop pairs. However, for overlap-based loop pairs, LCDNet’s performance declined, while our method still maintains superior registration capabilities, demonstrating its robustness for viewpoint changes. In terms of alignment accuracy, our method also significantly outperforms other baselines. We also present the qualitative results in Fig. 3. As can be seen, our method achieves better alignment even in low overlap loop closure, attributing to the semantically-guided coarse-fine-refine point cloud alignment process. Tab. VI shows the results on the KITTI-360 dataset with distance-based close loop. It can be observed that LCDNet exhibits the best generalization ability, but our method remains competitive and has the best alignment accuracy, considering both RTE and RRE. We also report the average running time of pose estimation on our machine. From the comparison, our method boasts the fastest running speed. The total runtime of our system, incorporating semantic segmentation, loop closure detection and poses estimation, is 98.1 ms, less than the data acquisition time of 100 ms for a typical rotational LiDAR sensor, thereby rendering it suitable for real-time SLAM systems.
Figure 3: The qualitative comparison of loop pose estimation on the KITTI dataset using overlap-based loop pairs. Dashed ellipses are directly annotated on the registration results, while solid boxes indicate local magnification. The top-left corner displays the ground truth distance and yaw angle difference of the loop pair.
IV-D Performance on SLAM System
We integrate SGLC into A-LOAM odometry to eliminate cumulative errors for evaluating the accuracy of SLAM trajectories, and compare with baseline integrated with BoW3D. We utilize a Incremental Smoothing and Mapping (iSAM2) [38] based pose-graph optimization (PGO) [^2] https://github.com/gisbi-kim/SC-A-LOAM framework to build a factor graph for managing the the keyframe poses. If a loop closure is detected, we will add this loop closure constraint into the factor graph to eliminate errors and ensure global trajectories consistency. As shown in Fig. 4, we present the trajectory error analysis between keyframe and ground truth poses. From the results, our method significantly enhances the SLAM system by reducing the cumulative odometry error, and shows superior performance compared to the baseline that incorporates BoW3D. This improvement is attributed to the excellent loop closing capability of SGLC.
@@ IV-E Ablation Study
In this section, we conduct a series of ablation studies on KITTI sequence 00 and 08 and report the mean metrics to evaluate the effectiveness of designed components, and analyze their runtime. The results are presented in Tab. VII. [[[[A]]]] is our baseline by removing those components listed in Tab. VII. Comparing [[[[A]]]] and [[[[B]]]], it is demonstrated that incorporating the global properties of instance nodes can enhance the performance of loop closing. As anticipated, it increases each node’s distinctiveness, which is beneficial for finding the correct node correspondences. For [[[[C]]]], we can see the effectiveness of outlier pruning in improving registration accuracy with the same number of RANSAC iterations as [[[[B]]]]. In experiment [[[[D]]]], [[[[E]]]] and [[[[F]]]], dense points registration and point-to-plane alignment are applied mainly for loop scan determined after geometric verification, so they do not affect the results of loop closure detection. The results clearly show that both dense points registration and point-to-plane alignment can improve registration performance, and we can get optimal results only when they operate in conjunction. The comparison of [[[[F]]]], [[[[G]]]], [[[[H]]]] also indicates that even direct point-to-point ICP or point-to-plane alignment on the raw scan cannot get optimal results and significantly increase computational costs on direct point-to-point ICP, which demonstrates that our registration process effectively utilizes foreground and background. Besides, each component in our framework exhibits high execution efficiency.
TABLE VII: Ablation studies. “Node Glo.” denotes the global properties of instance nodes. “Out. Pru.” denotes outlier pruning for node matches. “Den. ICP” denotes dense points registration, i.e., $\mathbf{T}_{\text{icp}}$. “P-P Ali.” denotes point-to-plane alignment, i.e., $\mathbf{T}_{\text{refine}}$.
Node Glo. Out. Pru. Den. ICP P-P Ali. F1max RR(%) RTE(m) RRE() Time (ms)
[[[[A]]]] 0.989 97.30 0.21 1.23 9.1
[[[[B]]]] 0.993 97.76 0.20 1.17 +0.5
[[[[C]]]] 0.994 98.29 0.18 0.99 +1.2
[[[[D]]]] - 99.98 0.07 0.38 +11.8
[[[[E]]]] - 99.71 0.06 0.34 +8.5
[[[[F]]]] - 100 0.06 0.31 +20.2
[[[[G]]]] $\bullet$ - 100 0.06 0.33 +86.1
[[[[H]]]] $\bullet$ - 99.65 0.06 0.34 +9.4
  • + denotes the time increment compared to [[[[A]]]]. $\bullet$ denotes direct point-to-point ICP or point-to-plane alignment for the raw scan.